OpenStudy (asiannerd):

Solve cos(x)(cos(x)-1)=0

OpenStudy (dessyj1):

You should try and expand the right side.

OpenStudy (dessyj1):

left* side

OpenStudy (dessyj1):

What have you tried doing?

OpenStudy (sshayer):

\[either~\cos x=0=\cos \left( n+1 \right)\frac{ \pi }{ 2 },x=\left( n+1 \right)\frac{ \pi }{ 2 }\] where n is an integer. \[or~\cos x=-1=\cos \left( 2n+1 \right)\pi,x=\left( 2n+1 \right)\pi\] where n is an integer.

OpenStudy (sshayer):

correction \[\cos x-1=0,\cos x=1=\cos 2n \pi,x=2n \pi\] where n is an integer.